if Log2=.3010 and log3=.4771, then find
(A) log1/25 (b) log3.002

Answer by Alan3354(69443)

if Log2=.3010 and log3=.4771, then find
(A) log1/25
----------------
1/25 = log(0.04) = log(2^2/100)
= 2log(2) - log(100)
= 0.602 - 2
= -1.398
===============
(b) log3.002
3.002 can't be expressed as a product of 2, 3 & 10